Running on a treadmill is slightly easier than running outside because there is no drag force to work against. Suppose a 60 kg runner completes a 5.0 km race in 23 minutes. the drag force on the runner during the race is 6.8 N when the cross section area is .72m^3.

Part B

What is this force as a fraction of the runner's weight?
Answer #1

Runners weight = W = mg = 60 × 9.8 = 588 N

Drag force = F = 6.8 N

F/W = 6.8/588 = 0.01156